OpenStudy (anonymous):

Adding and simplifying the fractions: 3/8+1/2+2/3.. I just got into this class and i don't understand how to do this..

OpenStudy (amistre64):

you need to fit them all into something that they can relate to .... or algebra thru it

OpenStudy (anonymous):

How so?

OpenStudy (amistre64):

\[\frac{3}{8}+\frac{1}{2}+\frac{2}{3}=N\] \[\frac{3(8)}{8}+\frac{1(8)}{2}+\frac{2(8)}{3}=(8)N\] \[3+4+\frac{16}{3}=8N\] \[3(3)+4(3)+\frac{16(3)}{3}=8(3)N\] \[9+12+16=24N\] \[9+28=24N\] \[37=24N\] \[\frac{37}{24}=N\]

OpenStudy (amistre64):

do you see what i did?

OpenStudy (amistre64):

if it doesnt make sense, youll have to speak up, and if it dies make sense ... you can still speak up :)

OpenStudy (anonymous):

wow, thanks alot

OpenStudy (anonymous):

that seems overly complicated unless I'm missing something. Wouldn't you just create common denominators?

OpenStudy (amistre64):

its the same process; just in a difference light

OpenStudy (anonymous):

its still complicated but atleast not i can grasp onto the concept of it. I will be able to look back at what you did to understand it later , so thanks alot

OpenStudy (amistre64):

all your doing is assigning it to some arbitrary value "N" and solveing for N

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

No, I like how you did it. It just seems like a lot of arbitrary steps as opposed to just multiplying it out. I dunno, I haven't taken a fraction class in like 10 years.

OpenStudy (anonymous):

3/8+1/2+2/3 =(3*3+4*12+2*8)/24 =9+48+16/24 =73/24

OpenStudy (amistre64):

\[\frac{3}{8}+\frac{1}{2}+\frac{2}{3}=N;\ cd,\ 24?\] \[\frac{3}{8}\frac{3}{3}+\frac{1}{2}\frac{12}{12}+\frac{2}{3}\frac{8}{8}=N\] \[\frac{9}{24}+\frac{12}{24}+\frac{16}{24}=N\] \[\frac{9+12+16}{24}=N\] \[\frac{9+28}{24}=N\] \[\frac{37}{24}=N\]

OpenStudy (anonymous):

Yeah, when you write it out I guess it's about the same. Don't really think about it much when you do it in your head.

OpenStudy (amistre64):

:) i like to think of math as being fluid and bendable; as opposed to a strict formula that has to be done by the "teachers" methods ... just my thoughts

OpenStudy (amistre64):

the hard part with fractions is trying to determine a suitable common denominator, so the first method just does that for you

OpenStudy (anonymous):

Yeah, that makes sense when you start getting into larger numbered fractions.